5.You are 4 times as old as your sister. In 8 years time you will be twice as old as your sister. What are your ages now?
Please explains also why u do not add 8 to 4x during the equation and why does this have to have two equations

Answer by ankor@dixie-net.com(22740)

Let y = your age present age
Let s = sister's present age
:
write an equation for each statement
:
You are 4 times as old as your sister.
y = 4s
;
In 8 years time you will be twice as old as your sister.
y + 8 = 2(s+8)
y + 8 = 2s + 16
y = 2s + 16 - 8
y = 2s + 8
Replace y with 4s
4s = 2s + 8
4s - 2s = 8
2s = 8
s = 4 yrs is sister's age
and
4(4) = 16 yrs is your age
:
:
Please explains also why u do not add 8 to 4x during the equation
"look at our equations, you see we subtracted 8 and got y = 2s+8"
and why does this have to have two equations
"When you have two unknowns, two equations are required
you skipped the first one, since you know y = 4s"